Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
boldrini, casanueva, bersanetti, pinto - 17:37 Tuesday 17 February 2026 (68703) Print this report
SRCL_SET stability range exploration

After the maintenance we set the ITF in LN2 and allowed about 10 minutes for the state to stabilize, before proceeding with the exploration of the stability range of SRCL_SET.

We closed BS_TX in full bandwidth and opened the SRCL_SET servo, then started changing SRCL_SET in steps of 10, one minute per ramp, allowing 20 minutes for each new state to stabilize and gather data. We managed to make a three points scan:

name final UTC DCP SRCL_SET SRCL_SET_TOT
reference 13:26:35 415 Hz 0 25.6
step 1 13:52:18 412 Hz -10 17.5
step 2 14:14:00 403 -20 7.5
step 3 \ 380 -30 -2.5

the ITF unlocked a few minutes after the last change of SRCL_SET (Fig.1).

We relocked the interferometer and decided to explore the margin of the scan more granularly. After reaching LN2 we set BS_TX in fbw and disabled the SRCL_SET servo again, then we set SRCL_SET to -20 with a two minutes ramp, allowed about 15 minutes to stabilize. We started changing the SRCL_SET in steps of 2, with one minutes long ramps, but realized after the first step that very little changes with such a small step while the duration of the test becomes considerable. For the next steps, we changed SRCL_SET in by 4, with two minutes long ramps.

name final UTC DCP SRCL_SET SRCL_SET_TOT
reference 15:08:00 425 Hz 0 32.0
step 0 15:24:30 390 Hz -20 12.0
step 1 15:41:50 390 Hz -22 10.0
step 2 17:00:28 387 Hz -26 6.0
step 3 17:18:40 382 Hz -30 2
step 4 \ 380 Hz -34 -2

The ITF unlocked about 10 minutes after step 4 (Fig.2)

The sensitivity was unexpectdly high for LN2, but the correct reconstruction of Hrec needs to be checked.

Images attached to this report
Comments to this report:
bersanetti, rolland, mwas - 18:32 Tuesday 17 February 2026 (68705) Print this report

The issue could be related to the fact that one of the relays of the NE actuator (REL21) was not correctly set to the LowNoise2 mode in the second lock. If the GUI shows the relays in the standard ordering, this should mean that the DR coil was affected, and this confirms the checks done offline on the Hrec reconstruction and the NE TF monitoring in particular.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×